# The DREAMS Project: Disentangling the Impact of Halo-to-Halo Variance and Baryonic Feedback on Milky Way Dark Matter Density Profiles

### v1

@@ 2.2 Reconstruction of Dark Matter Density
We sample radially in $300$ bins ranging from $2.8$ times the softening length ($\sim 1.2$ kpc) to $600$ kpc. [^3] For distances greater than $\sim 2.8\times$ the softening length, gravity is fully Newtonian, whereas below this scale, the shape of the potential is impacted significantly by the choice of softening length (Ludlow et al., 2020). We therefore omit any radial bins at <1.2<1.2 kpc from the center of the halo. At each radius, we sample this reconstructed density field uniformly along a spherical shell with $10^{3}$ points. We take the average density on this shell as the density at that particular radial location. We also calculate a standard deviation of densities at each radial bin as a measure of azimuthal variations of the density within the shell. The typical standard deviation of densities is $\lesssim 0.1$ dex at all radii within $R_{200}$, suggesting that the halos are reasonably spherical—even while including the contribution from satellites. We leave a more careful examination of the three-dimensional shape of the dark matter halos for a future work.
2.2.1 Fitting Analytic Profiles
In Section 3.2.2, we fit each profile with an analytic function to get a more quantitative understanding of the shape of the dark matter density profiles. We fit the profiles with a generalized NFW (gNFW) profile of the form
@@ 2.2.1 Fitting Analytic Profiles
where $\rho_{s}$ and $r_{s}$ are the scale density and radius (respectively), $\gamma$ is the inner slope, $\beta$ is the outer slope, and $\alpha$ regulates how sharply the transition between the two occurs (e.g., Jaffe, 1983; Hernquist, 1990; Zhao, 1996; Merritt et al., 2006; Di Cintio et al., 2014). In the case that $\alpha=1,\ \beta=3,~{\rm and}~\gamma=1$ the gNFW profile is identical to that of the canonical NFW profile (i.e., Equation 1).
@@
where $\log\rho$ is the measured average density profile in each radial bin, $\log\rho_{\rm model}(r|\phi)$ is the gNFW prediction at radius $r$ given fit parameters $\phi$, and $\sigma_{\rho}$ is the standard deviation of densities about the average profile. We adopt broad Gaussian priors centered on the values obtained from the previous steps using a factor of $5\times$ the uncertainty on the non-linear least squares fits (given by square root of covariance matrix along the diagonal) as the width of the Gaussian prior. This ensures the gNFW fits are not restricted to NFW-like solutions but instead use the NFW fit only to guide initialization. Finally, we limit this fitting procedure to radii less than $R_{200}$.
2.3 Emulators
Throughout this work, we employ a neural network emulator to act as a multi-dimensional interpolator of the parameter space sampled in the DREAMS simulations. These emulators are trained to predict summary statistics of the simulations as a function of the underlying cosmological and astrophysical parameters. By learning the complex, non-linear mappings between input parameters and simulated outputs, the emulator efficiently explores parameter dependencies and makes predictions without the computational cost of running additional simulations. The emulator also provides estimates of prediction uncertainty, enabling robust marginalization over multiple parameters at once. We briefly describe our emulators in this section, noting that they are very similar in both spirit and implementation to that of Rose et al. (2025a).
Our emulators take the five simulation parameters as input ($\bar{e}_{w}$, $\kappa_{w}$, $\epsilon_{f,\,\mathrm{high}}$, $\Omega_{\rm M}$, and $\sigma_{8}$) as well the mass of the halo ($M_{\rm halo}$). These inputs are fed into a series of fully connected layers which output a prediction for the mean and standard deviation of the quantity of interest. The loss function is defined as
@@ 2.3 Emulators
where $x$ is the true value of the target in each simulation, $\mu$ is the predicted mean, and $\sigma$ is the predicted standard deviation. This loss function minimizes the mean-squared-error of the data (first term) while also ensuring that the measured deviation is reproduced (second term). It allows the network to learn both accurate predictions and accurate mean and variance estimation (Jeffrey & Wandelt, 2020; Villaescusa-Navarro et al., 2022; Rose et al., 2025a).
We split the data into training, validation, and test sets, which comprise $80\%$, $10\%$, and $10\%$ of the dataset, respectively. We optimize our models using the optuna package (Akiba et al., 2019) to select: (i) the number of fully connected layers, (ii) the nodes in each layer, (iii) the learning rate, (iv) the weight decay, and (v) the dropout rate. Following directly from Rose et al. (2025a), we vary the number of layers from $1$ to $5$, the number of neurons from $4$ to $10^{3}$, the learning rate from $10^{-5}$ to $10^{-1}$, the weight decay from $10^{-8}$ to $1$, and the dropout rate from $0.2$ to $0.8$. The hyperparameter tuning is done using $50$ trials consisting of $500$ epochs each.
We employ emulators in six places in this work to output gNFW shape parameters (Figures 3 and 4 in Section 3.2), dark matter mass growth in the inner regions of halos (Figure 5 in Section 3.3), the stellar mass and black hole mass of the central galaxies (Figure 6 in Section 3.3), the lifetime feedback energy output by AGN (Figure C1 in Appendix C), and fitting parameters for our adiabatic contraction calculations (Figure D1 in Appendix D). In each case, we hyperparameter optimize and train ten instances of the emulator with different random initializations of the weights (for a total of $50$ models created for this work). In all the results showing predictions from the emulators, we report the mean of the predictions from the ten instances. Each instance we report in this work is comprised of emulation of $1000$ halos of a single parameter variation keeping the other parameters fixed at their fiducial TNG values and halo mass fixed to $M_{\rm Halo}=10^{12}~{\rm M}_{\odot}$. This means that, e.g., for each panel in Figure 3 there are a total of $10\times 1000$ draws from the emulators (for a total of 60,000 in the whole figure). Figure A1 provides an example validation of the emulator that predicts the gNFW shape parameters.
2.3.1 Definition of Halo-to-Halo Variation
@@
2.4 Halo Weighting Scheme
@@
Figure 2: Dark Matter Density Profiles in the DREAMS CDM Milky Way-Mass Suite. The large left-hand panel shows the dark matter density as a function of radius for all halos in the DREAMS CDM Milky Way-mass suite from $\sim 1.2$ kpc to $600$ kpc (normalized by $R_{200}$). The red line and shaded region corresponds to the weighted mean and plus/minus standard deviation of the individual profiles (see Section 2.4 and Rose et al. (2025b) for details on the weighting procedure). The values in the bottom left of the large panel correspond to the weighted standard deviation of densities at $0.01R_{200}$ ($\sim 2~{\rm kpc}$), $0.1R_{200}$ ($\sim 20$ kpc), and $R_{200}$ ($\sim 200$ kpc). The panels on the right show how the density depends on the upper (dotted orange) and lower (dashed blue) $10\%$ range for each of the parameters (compared to the average $10\%$ of the parameter ranges, $\rho_{\rm avg}$): $\bar{e}_{w}$ (top left), $\kappa_{w}$ (top middle), $\epsilon_{f,\,\mathrm{high}}$ (top right), $\Omega_{\rm M}$ (bottom left), and $\sigma_{8}$ (bottom middle). The colored line in each panel is the median density profile while the colored shaded region represents the $25^{\rm th}$ and $75^{\rm th}$ percentile. Overall, the profiles are self similar, with a modest amount of scatter. Moreover, the baryonic parameters have the strongest effect on the central densities.
To quantify how reasonable the DREAMS parameter variations are, we follow the halo weighting scheme outlined in Rose et al. (2025b). That work uses a “pseudo-posterior” to determine how well a simulated halo’s properties compare to well-established scaling relations from observations (in the case of Rose et al. 2025b, the stellar mass–halo mass relation; Wang et al. 2024). Notably, since the properties of individual halos depend on their detailed formation histories, the weighting is not calcuated on a halo-by-halo basis. Instead, we separate the model parameter space into $30$ evenly-space bins along each dimension ($\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,{\rm high}}$) [^4] The version of the weighting scheme presented here does not include variations of $\Omega_{\rm M}$ or $\sigma_{8}$. Empirically, we find that the weighting scheme does not put strong constraints on these two cosmological parameters, suggesting that they do not play a significant role in setting the stellar mass–halo mass relation. and apply the weights to halos that fall within each bin. Rose et al. (2025b) use a normalizing flow emulator (similar to that of Nguyen et al., 2024) to generate samples consistent with the underlying DREAMS data to more finely sample the high-dimensional parameter space. A sample’s residual given some set of input parameters, $\theta$, is calculated such that
@@
where $\tau$ is a free parameter set to $0.2~\log[{\rm M}_{\odot}]$ (see Rose et al., 2025b). We note that the results of this work do not depend sensitively on the choice of $\tau$. Finally, we normalize the weights such that
@@
The results of this weighting procedure can be seen in Figure 1 of Rose et al. (2025b). To briefly summarize, we down-weight the extreme values of $\bar{e}_{w}$ and $\kappa_{w}$ and the lower-than-fiducial values of $\epsilon_{f,\,\mathrm{high}}$, but do not place strong constraints on $\Omega_{\rm M}$ or $\sigma_{8}$.
We note that our main conclusions are not particularly sensitive to using this weighting scheme (see Appendix B). However, since there are halos with unrealistically large/small stellar masses given the halo mass of the Milky Way, we aim to down weight their contributions.
3 Results
3.1 Dark Matter Density Profiles
Figure 2 shows the dark matter density profiles for all of the DREAMS halos (large panel on left). To compare the results across simulations, the radii are normalized by $R_{200}$ for each galaxy. The weighted mean and standard deviation are shown by the red line and shaded band, respectively; these correspond to the density profiles of those Milky Way halos whose central galaxies are most consistent with the observed stellar mass–halo mass relation.
On the whole, the dark matter profiles are largely consistent in terms of normalization, which is, in part, by construction of the selection function of the DREAMS halos. There is a tight (weighted) spread in densities of $0.13$ dex at $0.1R_{200}$ ($\sim 20~{\rm kpc}$). At larger radii, the spread increases by a factor of $\sim 2$ ($0.31~{\rm dex}$ at $R_{200}$) which is a feature of the halos containing many subhalos at these large radii. Note that the presence of more satellites leads to occasional “spikes” at large radii. In the inner-most regions of the halo, $0.01R_{200}$ ($\sim 2~{\rm kpc}$), there is a much wider distribution of densities with a spread of $0.31~{\rm dex}$ (a factor of $\gtrsim 2$ increase from $0.1R_{200}$). Interestingly, this trend is consistent with massive halos with mass above $10^{13}~{\rm M}_{\odot}$ (see left panel of Figure 2 in Farahi et al., 2022).
@@ 3.1 Dark Matter Density Profiles
The lower $\bar{e}_{w}$ values correspond to a median density ratio of $1.278_{-0.315}^{+0.185}$ ($\sim 30\%$ increase), while the upper $\bar{e}_{w}$ values have $0.737_{-0.281}^{0.213}$ ($\sim 30\%$ decrease). However, within the spread, the lower $\bar{e}_{w}$ variations are consistent with no significant change in the inner density. For $\kappa_{w}$, the lower values correspond to a ratio of $1.697^{+0.508}_{-0.463}$ ($\sim 70\%$ increase), and the upper values correspond to a ratio of $0.737^{+0.213}_{-0.281}$ ($\sim 30\%$ decrease). Notably, neither the upper nor lower $\kappa_{w}$ ranges are consistent with the average profile, at least within the inner-quartile range. Interestingly, the enhancement/decrease caused by $\kappa_{w}$ (which controls wind speed in our simulations) extends to much larger radii than those of $\bar{e}_{w}$ (which controls wind energy).
@@
In terms of $\epsilon_{f,\,\mathrm{high}}$, the upper values are consistent with the average profile (ratio of $1.045_{-0.486}^{+0.387}$), while the lower values are very slightly below the average profile (ratio of $0.782_{-0.295}^{0.207}$). While very slight, this decrease in central density with decreased $\epsilon_{f,\,\mathrm{high}}$ is likely related to a self-regulation of the black holes which we discuss in more detail in Section 3.3.
For both the $\Omega_{\rm M}$ and $\sigma_{8}$ variations, both the lower and upper ranges are consistent with the average profile.
3.2 Analytic Fits to Profiles
@@
Figure 3: Dependence of generalized NFW scale density ($\rho_{s}$) on Astrophysics, Cosmology, and Halo Mass. Predictions from an ensemble of emulators (see Section 2.3) for the scale density, $\rho_{s}$, of the DREAMS CDM Milky Way-mass density profiles. We show the predictions made by the emulator as a function of $\bar{e}_{w}$ (top left), $\kappa_{w}$ (top middle), $\epsilon_{f,\,\mathrm{high}}$ (top right), $\Omega_{\rm M}$ (bottom left), $\sigma_{8}$ (bottom middle), and $M_{\rm halo}$ (bottom right). The shaded regions represent the one standard deviation uncertainty of the predictions based on the individual model prediction uncertainty as well as variance across the emulators (via Equation 12). The vertical red solid line at the top of each panel corresponds to the fiducial TNG value (see Table 1). In every panel, the non-varied parameters are fixed to their fiducial values and $M_{\rm Halo}=10^{12}~{\rm M}_{\odot}$ unless explicitly varied. While not shown, the emulator also predicts the scale radius. The trends for the scale radius are identical to those shown here, just in the opposite direction (i.e., increasing $\kappa_{w}$ increases $r_{s}$). We find that only $\kappa_{w}$ drives variation in $\rho_{s}$ more significant than the halo-to-halo variation.
@@
Figure 4: Dependence of generalized NFW shape parameters on $\mathbf{\bar{e}_{w}}$. Predictions from our neural network emulator for the dependence of the inner slope ($\gamma$; dot-dashed line), transition rate ($\alpha$; dashed line), and outer slope ($\beta$) of the best-fit gNFW profiles on the supernova wind energy ($\bar{e}_{w}$). As a point of reference, the dotted gray lines show the canonical NFW profile values ($\alpha=\gamma=1$ and $\beta=3$). The short vertical solid line at the top corresponds to the fiducial TNG value of $\bar{e}_{w}=3.6$ (see Table 1; Section 2.1.1). The average prediction lines are the mean of our ten emulators while the shaded regions approximate the halo-to-halo variation (although there is also a small contribution from the combination of emulators, see Section 2.3 for details). We note that, within the uncertainty due to halo-to-halo variation, the shape parameters have no strong dependence on any of the other DREAMS simulation parameters or halo mass.
3.2.1 Scale Density

### v2

@@ II.2 Reconstruction of Dark Matter Density
We sample radially in $300$ bins ranging from $2.8$ times the softening length ($\sim 1.2$ kpc) to $600$ kpc. [^3] For distances greater than $\sim 2.8\times$ the softening length, gravity is fully Newtonian, whereas below this scale, the shape of the potential is impacted significantly by the choice of softening length. Although previous work (e.g., Ludlow et al., 2020) suggests that full numerical convergence of dark matter density profiles may require radii somewhat larger than $2.8\times$ the softening, our analysis does not rely on the detailed structure at the smallest resolved radii. In practice, many of the dominant differences between hydrodynamical and DMO runs extend to radii closer to $\sim 0.02R_{\rm 200}$ ($\sim 5$ kpc or $\sim 12\times$ the softening length; see, e.g., Figures 27, and 8) At each radius, we sample this reconstructed density field uniformly along a spherical shell with $10^{3}$ points. We take the average density on this shell as the density at that particular radial location. We also calculate a standard deviation of densities at each radial bin as a measure of azimuthal variations of the density within the shell. The typical standard deviation of densities is $\lesssim 0.1$ dex at all radii within $R_{200}$, suggesting that the halos are reasonably spherical—even while including the contribution from satellites. We leave a more careful examination of the three-dimensional shape of the dark matter halos for future work.
II.2.1 Fitting Analytic Profiles
In Section III.2.2, we fit each profile with an analytic function to get a more quantitative understanding of the shape of the dark matter density profiles. We fit the profiles with a generalized NFW (gNFW) profile of the form
@@ II.2.1 Fitting Analytic Profiles
where $\rho_{s}$ and $r_{s}$ are the scale density and radius (respectively), $\gamma$ is the inner slope, $\beta$ is the outer slope, and $\alpha$ regulates how sharply the transition between the two occurs (e.g., Jaffe, 1983; Hernquist, 1990; Zhao, 1996; Merritt et al., 2006; Di Cintio et al., 2014). In the case that $\alpha=1,\ \beta=3,~{\rm and}~\gamma=1$, the gNFW profile is identical to that of the canonical NFW profile (i.e., Equation 1).
@@
where $\log\rho$ is the measured average density profile in each radial bin, $\log\rho_{\rm model}(r|\phi)$ is the gNFW prediction at radius $r$ given fit parameters $\phi$, and $\sigma_{\rho}$ is the standard deviation of densities about the average profile. We adopt broad Gaussian priors centered on the values obtained from the previous steps using a factor of $5\times$ the uncertainty on the non-linear least squares fits (given by the square root of the covariance matrix along the diagonal) as the width of the Gaussian prior. This ensures the gNFW fits are not restricted to NFW-like solutions but instead use the NFW fit only to guide initialization. Finally, we limit this fitting procedure to radii less than $R_{200}$.
II.3 Emulators
Throughout this work, we employ a neural network emulator to act as a multi-dimensional interpolator of the parameter space sampled in the DREAMS simulations. These emulators are trained to predict summary statistics of the simulations as a function of the underlying cosmological and astrophysical parameters. By learning the complex, non-linear mappings between input parameters and simulated outputs, the emulator efficiently explores parameter dependencies and makes predictions without the computational cost of running additional simulations. The emulator also provides estimates of prediction uncertainty, enabling robust marginalization over multiple parameters at once. We briefly describe our emulators in this section, noting that they are very similar in both spirit and implementation to that of R25a.
Our emulators take the five simulation parameters as input ($\bar{e}_{w}$, $\kappa_{w}$, $\epsilon_{f,\,\mathrm{high}}$, $\Omega_{\rm M}$, and $\sigma_{8}$) as well the mass of the halo ($M_{\rm halo}$). These inputs are fed into a series of fully connected layers, which output a prediction for the mean and standard deviation of the quantity of interest. The loss function is defined as
@@ II.3 Emulators
where $x$ is the true value of the target in each simulation, $\mu$ is the predicted mean, and $\sigma$ is the predicted standard deviation. This loss function minimizes the mean-squared-error of the data (first term) while also ensuring that the measured deviation is reproduced (second term). It allows the network to learn both accurate predictions and accurate mean and variance estimation (Jeffrey & Wandelt, 2020; Villaescusa-Navarro et al., 2022, R25a).
We split the data into training, validation, and test sets, which comprise $80\%$, $10\%$, and $10\%$ of the dataset, respectively. We optimize our models using the optuna package (Akiba et al., 2019) to select: (i) the number of fully connected layers, (ii) the nodes in each layer, (iii) the learning rate, (iv) the weight decay, and (v) the dropout rate. Following R25a, we vary the number of layers from $1$ to $5$, the number of neurons from $4$ to $10^{3}$, the learning rate from $10^{-5}$ to $10^{-1}$, the weight decay from $10^{-8}$ to $1$, and the dropout rate from $0.2$ to $0.8$. The hyperparameter tuning is done using $50$ trials consisting of $500$ epochs each.
We employ emulators to output gNFW shape parameters (Figures 3 and 4 in Section III.2), dark matter mass growth in the inner regions of halos (Figure 5 in Section III.3), the stellar mass and black hole mass of the central galaxies (Figure 6 in Section III.3), the lifetime feedback energy output by AGN (Figure C1 in Appendix B), and fitting parameters for our adiabatic contraction calculations (left panel of Figure D1 in Appendix C). In each case, we hyperparameter optimize and train ten instances of the emulator with different random initializations of the weights (for a total of $50$ models created for this work). In all the results showing predictions from the emulators, we report the mean of the predictions from the ten instances. Each instance we report in this work is comprised of emulation of $1000$ halos of a single parameter variation keeping the other parameters fixed at their fiducial TNG values and halo mass fixed to $M_{\rm Halo}=10^{12}~{\rm M}_{\odot}$. This means that, e.g., for each panel in Figure 3 there are a total of $10\times 1000$ draws from the emulators (for a total of 60,000 in the whole figure). Figure A1 provides an example validation of the emulator that predicts the gNFW shape parameters.
II.3.1 Definition of Halo-to-Halo Variation
@@
II.4 Halo Weighting Scheme
@@
Figure 2: Dark Matter Density Profiles in the DREAMS CDM Milky Way-Mass Suite. The large left-hand panel shows the dark matter density as a function of radius for all halos in the DREAMS CDM Milky Way-mass suite from $\sim 1.2$ kpc to $600$ kpc (normalized by $R_{200}$). The brown line and shaded region correspond to the weighted mean and plus/minus standard deviation of the individual profiles (see Section II.4 and R25b for details on the weighting procedure). The values in the bottom left of the large panel correspond to the weighted standard deviation of densities at $0.01R_{200}$ ($\sim 2~{\rm kpc}$), $0.1R_{200}$ ($\sim 20$ kpc), and $R_{200}$ ($\sim 200$ kpc). The panels on the right show how the density depends on the upper (dotted orange) and lower (dashed blue) $10\%$ range for each of the parameters (compared to the average $10\%$ of the parameter ranges, $\rho_{\rm avg}$): $\bar{e}_{w}$ (top left), $\kappa_{w}$ (top middle), $\epsilon_{f,\,\mathrm{high}}$ (top right), $\Omega_{\rm M}$ (bottom left), and $\sigma_{8}$ (bottom middle). The colored line in each panel is the median density profile, while the colored shaded region represents the $25^{\rm th}$ and $75^{\rm th}$ percentile. Overall, the profiles are self-similar, with a modest amount of scatter. Moreover, the baryonic parameters have the strongest effect on the central densities.
To quantify how reasonable the DREAMS parameter variations are, we follow the halo weighting scheme outlined in R25b. That work uses a “pseudo-posterior” to determine how well a simulated halo’s properties compare to well-established scaling relations from observations (in the case of R25b, the stellar mass–halo mass relation; Wang et al. 2024). Notably, since the properties of individual halos depend on their detailed formation histories, the weighting is not calculated on a halo-by-halo basis. Instead, we separate the model parameter space into $30$ evenly-space bins along each dimension ($\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,{\rm high}}$) [^4] The version of the weighting scheme presented here does not include variations of $\Omega_{\rm M}$ or $\sigma_{8}$. Empirically, we find that the weighting scheme does not put strong constraints on these two cosmological parameters, suggesting that they do not play a significant role in setting the stellar mass–halo mass relation. and apply the weights to halos that fall within each bin. R25b use a normalizing flow emulator (similar to that of Nguyen et al., 2024) to generate samples consistent with the underlying DREAMS data to more finely sample the high-dimensional parameter space. A sample’s residual given some set of input parameters, $\theta$, is calculated such that
@@
where $\tau$ is a free parameter set to $0.2~\log[{\rm M}_{\odot}]$ (see R25b). We note that the results of this work do not depend sensitively on the choice of $\tau$. Finally, we normalize the weights such that
@@
The normalized weights admit a natural interpretation as importance weights (Farahi et al., 2026). The results of this weighting procedure can be seen in Figure 1 of R25b. To briefly summarize, we down-weight the extreme values of $\bar{e}_{w}$ and $\kappa_{w}$ and the lower-than-fiducial values of $\epsilon_{f,\,\mathrm{high}}$, but do not place strong constraints on $\Omega_{\rm M}$ or $\sigma_{8}$.
We note that our main conclusions are not particularly sensitive to using this weighting scheme since, as we will show, the density profiles themselves are not strongly dependent on the supernova and AGN feedback parameters (see discussion in Section III.1). However, since there are halos with unrealistically large/small stellar masses given the halo mass of the Milky Way, we aim to down weight their contributions.
III Results
III.1 Dark Matter Density Profiles
Figure 2 shows the dark matter density profiles for all of the DREAMS halos (large panel on left). To compare the results across simulations, the radii are normalized by $R_{200}$ for each galaxy. The weighted mean and standard deviation are shown by the brown line and shaded band, respectively; these correspond to the density profiles of those Milky Way halos whose central galaxies are most consistent with the observed stellar mass–halo mass relation.
On the whole, the dark matter profiles are largely consistent in terms of normalization, which is, in part, by construction of the selection function of the DREAMS halos. There is a tight (weighted) spread in densities of $0.13$ dex at $0.1R_{200}$ ($\sim 20~{\rm kpc}$). At larger radii, the spread increases by a factor of $\sim 1.5$ (increasing from $0.13~{\rm dex}$ to $0.31~{\rm dex}$ at $R_{200}$) which is a feature of the halos containing many subhalos at these large radii. Note that the presence of more satellites leads to occasional “spikes” at large radii. In the inner-most regions of the halo, $0.01R_{200}$ ($\sim 2~{\rm kpc}$), there is a much wider distribution of densities with a spread of $0.31~{\rm dex}$ (a factor of $\gtrsim 2$ increase from $0.1R_{200}$). Interestingly, this trend is consistent with massive halos with mass above $10^{13}~{\rm M}_{\odot}$ (see left panel of Figure 2 in Farahi et al., 2022).
@@ III.1 Dark Matter Density Profiles
At the inner most radii ($0.06R_{200}$), the lower $\bar{e}_{w}$ values correspond to a median density ratio of $1.278_{-0.315}^{+0.185}$ ($\sim 30\%$ increase), while the upper $\bar{e}_{w}$ values have $0.737_{-0.281}^{0.213}$ ($\sim 30\%$ decrease). However, within the spread, the lower $\bar{e}_{w}$ variations are consistent with no significant change in the inner density. For $\kappa_{w}$, the lower values correspond to a ratio of $1.697^{+0.508}_{-0.463}$ ($\sim 70\%$ increase), and the upper values correspond to a ratio of $0.737^{+0.213}_{-0.281}$ ($\sim 30\%$ decrease). Notably, neither the upper nor lower $\kappa_{w}$ ranges are consistent with the average profile, at least within the inner-quartile range. Interestingly, the enhancement/decrease caused by $\kappa_{w}$ (which controls wind speed in our simulations) extends to much larger radii than those of $\bar{e}_{w}$ (which controls wind energy).
@@
In terms of $\epsilon_{f,\,\mathrm{high}}$, the upper values are consistent with the average profile (ratio of $1.045_{-0.486}^{+0.387}$), while the lower values are very slightly below the average profile (ratio of $0.782_{-0.295}^{0.207}$). While very slight, this decrease in central density with decreased $\epsilon_{f,\,\mathrm{high}}$ is likely related to a self-regulation of the black holes which we discuss in more detail in Section III.3.
For both the $\Omega_{\rm M}$ and $\sigma_{8}$ variations, both the lower and upper ranges are consistent with the average profile within the uncertainty on the median. It is worth appreciating that if we were to consider wider variations on $\Omega_{\rm M}$ and $\sigma_{8}$ as in the CAMELS simulations (Villaescusa-Navarro et al. 2021; $\Omega_{\rm M}\in[0.1,0.5]$ and $\sigma_{8}\in[0.6,1.0]$) it is possible that deviations from the average profile might be more significant.
III.2 Analytic Fits to Profiles
@@
Figure 3: Dependence of generalized NFW scale density ($\rho_{s}$) on Astrophysics, Cosmology, and Halo Mass. Predictions from an ensemble of emulators (see Section II.3) for the scale density, $\rho_{s}$, of the DREAMS CDM Milky Way-mass density profiles. We show the predictions made by the emulator as a function of $\bar{e}_{w}$ (top left), $\kappa_{w}$ (top middle), $\epsilon_{f,\,\mathrm{high}}$ (top right), $\Omega_{\rm M}$ (bottom left), $\sigma_{8}$ (bottom middle), and $M_{\rm halo}$ (bottom right). The shaded regions represent the one standard deviation uncertainty of the predictions based on the individual model prediction uncertainty as well as variance across the emulators (via Equation 12). The vertical red solid line at the top of each panel corresponds to the fiducial TNG value (see Table 1). In every panel, the non-varied parameters are fixed to their fiducial values and $M_{\rm Halo}=10^{12}~{\rm M}_{\odot}$ unless explicitly varied. While not shown, the emulator also predicts the scale radius. The trends for the scale radius are identical to those shown here, just in the opposite direction (i.e., increasing $\kappa_{w}$ increases $r_{s}$). We find that only $\kappa_{w}$ drives variation in $\rho_{s}$ more significant than the halo-to-halo variation.
@@
Figure 4: Dependence of generalized NFW shape parameters on $\mathbf{\bar{e}_{w}}$. Predictions from our neural network emulator for the dependence of the inner slope ($\gamma$; dot-dashed line), transition rate ($\alpha$; dashed line), and outer slope ($\beta$) of the best-fit gNFW profiles on the supernova wind energy ($\bar{e}_{w}$). As a point of reference, the dotted gray lines show the canonical NFW profile values ($\alpha=\gamma=1$ and $\beta=3$). The short vertical solid line at the top corresponds to the fiducial TNG value of $\bar{e}_{w}=3.6$ (see Table 1; Section II.1.1). The average prediction lines are the mean of our ten emulators while the shaded regions approximate the halo-to-halo variation (although there is also a small contribution from the combination of emulators, see Section II.3 for details). We note that, within the uncertainty due to halo-to-halo variation, the shape parameters have no strong dependence on any of the other DREAMS simulation parameters or halo mass.
III.2.1 Scale Density